Parameter = 46, L= 3*W - 5
Found 2 solutions by checkley77, richwmiller:
Answer by checkley77(12844)

46=2L+2W
46=2(3W-5)+2W
46=6W-10+2W
46=8W-10
8W=46+10
8W=56
W=56/8
W=7 ans. for the width.
L=3*7-5
L=21-5
L=16 ANS. FOR THE LENGTH.
PROOF:
46=2*16+2*7
46=32+14
46=46

Answer by richwmiller(17219)

So do you have a question about parameters or perimeters? They are not the same!